How many molecules of Ag2So4 will be made if 2.7 grams of AgNO3 are reacted?
AgNO3 + H2SO4 --> Ag2SO4 + HNO3
1
Expert's answer
2020-12-14T14:35:47-0500

The balanced equation for the reaction above is;

"2AgNO_3 + H_2SO_4 \to Ag_2SO_4 + 2HNO_3"


mass of AgNO3 = 2.7g

number of moles of AgNO3 = 2.7/169.87 = 0.016 moles


From the reaction above, 2 moles of AgNO3 will react to produce 1 mole of Ag2SO4

0.016 moles of AgNO3 will therefore produce 0.008 moles of Ag2SO4


0.008 moles = 0.008 × 6.02×10²³ = 4.82×10²¹ molecules


This means that 4.82×10²¹ molecules of Ag2SO4 would be made in the reaction.